IN A LOTTERY 10000 TICKETS ARE SOLD AND TEN PRIZES ARE AWARDED..WHAT IS THE PROBABILITY OF GETTING A PRIZE IF YOU BUY ONE TICKET..??????

Grade:12th pass

1 Answers

Sunil Raikwar
9 years ago
Here favourable outcomes=10
&,total outcome=1000
So probability wining a prize=10/1000=1/100
Thanks and Regards
Sunil Raikwar
askIITians faculty